OpenStudy (anonymous):

the expression 12/3+radical 3 is equivalent to

OpenStudy (asnaseer):

are you aware of this?\[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how would taht work?

OpenStudy (asnaseer):

it is commonly called the "difference of two squares" factorisation

OpenStudy (anonymous):

how would that plug in?

OpenStudy (asnaseer):

you can make use of that here as follows: what would you get if you multiplied \((3+\sqrt{3})\) by \((3-\sqrt{3})\) ?

OpenStudy (anonymous):

i did that and i got

OpenStudy (anonymous):

9 dont they cancel out?

OpenStudy (anonymous):

or would it be radical -9?

OpenStudy (asnaseer):